if a,b,c,d are positive real numbers such that a/3=a+b/4=a+b+c/5=a+b+c+d/6 then
a/b+2c+3d

Grade:

7 Answers

kasilaxmi
9 years ago
h
kasilaxmi
9 years ago
h
kasilaxmi
9 years ago
hi farzana
farzana
9 years ago
hello chandu.....
kasilaxmi
9 years ago
hey post raledu kada
Y RAJYALAKSHMI
9 years ago
Let a/3 = a + b / 4 = a + b + c / 5 = a + b + c + d/ 6 = k
then, a = 3k, b = k, c = k, d = k
 
a/b + 2c + 3d = 3k/k + 2k + 3k = 3k/6k = 1/2
Samrat Hadke
2 years ago
Correct option is
A

21

 

Given that, 3a=4a+b=5a+b+c=6a+b+c+d
 
So, we have 3a=4a+b
a=3b
 
We have, 3a=5a+b+c
a=3c
 
We have, 3a=6a+b+c+d
a=3d
 
So, the value of b+2c+3da=3a+32a+aa=2aa=21
 
Therefore, the correct option is A. ½